# Laypeople’s Egocentric Perceptions of Copyright for AI-Generated Art

### v1

@@ 3.4. Hypotheses
We build upon this prior work and study how egocentric effects may emerge in the context of GenAI. We formed two hypotheses concerning egocentric effects on lay perceptions of AI-generated images in relation to copyright. In H1, we compare the judgments of different participants about the same image. Namely, we hypothesize that the creator of an image will judge their image more favorably than other participants. In H2, we compare the judgments of the same participant about different images. We hypothesize that participants will judge images they create more favorably than they will judge images created by others.
H1)
@@
Egocentric effect between participants: Creators will 1) evaluate images more favorably and 2) be more likely to identify users as authors and right-holders when judging their own creations compared to invested and uninvested evaluators judging the same images.
H2)
@@
Egocentric effect between images: Invested evaluators will 1) evaluate images more favorably and 2) be more likely to identify users as authors and right-holders when judging their own creations than when judging other people’s images.
In our juried art exhibition, the highest scored submissions received rewards. Creators’ and invested evaluators’ outcomes depended on other participants’ assessments. Uninvested evaluators had no direct incentives to lie—the scores they provided had no impact on their own outcomes, since they did not take part in the competition. However, invested evaluators’ outcomes depended not only on the scores their submission received, but also on the scores other submissions received.
@@
Invested evaluators had the opportunity to inflate their own relative rating not only by giving high ratings to their own submission (in line with H2), but also by giving low scores to other people’s submissions. That is, they could benefit from sabotaging their competition. The problem of sabotage in competitions has been discussed extensively in economics: theory predicts its occurrence, and experiments show that sabotage is indeed empirically relevant (Chowdhury and Gürtler, 2015). Therefore, we hypothesize that the incentive structure associated with our juried art exhibition may lead those with something at stake (invested evaluators) to judge others’ submissions more harshly than those who do not have anything to gain from doing so (uninvested evaluators).
H3)
@@
Competition effect: Invested evaluators will 1) evaluate other people’s images less favorably and 2) be less likely to identify users as authors and right-holders than uninvested evaluators.
@@
Uninvested evaluators did not have an opportunity to interact with the GenAI model and to create AI-generated art in our experiments. On the other hand, creators and invested evaluators repeatedly interacted with the model throughout the tutorial and creation task, observing how changing their inputs influences the model’s outputs and learning how to use it to create AI-generated art. Research has identified several psychological phenomena, such as the mere-exposure effect (Zajonc, 1968; Montoya et al., 2017), practice and learning effects (Donovan and Radosevich, 1999), and the curse of knowledge (Hinds, 1999), which lead us to hypothesize that participants’ evaluations may be influenced by their interaction (or lack thereof) with the GenAI model. Consequently, we hypothesize that the experience of interacting with the GenAI model will influence participants’ perceptions about AI-generated art.
H4)
@@
Experience effect: Participants who interacted with the GenAI model (i.e., creators and invested evaluators) will 1) evaluate images and 2) attribute authorship and rights differently than those who did not use the GenAI model (i.e., uninvested evaluator).
@@
We emphasize that while we gather data on how participants attribute authorship and rights to various entities, our hypotheses focus on attributions of authorship and rights to one specific entity: users. We investigate the remaining entities exploratively.
@@ 3.5. Analysis Plan
We used regressions to analyze our data. We treated participants’ assessments of factors, authorship, and score evaluations as continuous dependent variables in linear regressions. Participants’ attributions of rights were treated as binary dependent variables and modeled using logit regressions. To account for repeated measurements across participants and images (i.e, participants evaluated several images, and images were evaluated by several participants), we initially planned to use mixed-effects linear regressions with crossed random intercepts for images and participants. However, due to convergence issues, we instead opted for regressions with two-way clustered standard errors (Correia et al., 2016).
Our primary independent variable is a dummy variable encoding both the i) treatment condition to which the participants were assigned and ii) whether the data point refers to an evaluation of their own image or someone else’s creation. Hence, our dummy variables has four levels representing 1) creators judging their own images; 2) invested evaluators rating their own creations; 3) invested evaluators assessing others’ images; and 4) uninvested evaluators judging others’ submissions.
We tested for differences between pairs of treatments by estimating their contrasts (i.e., by estimating the difference between the treatments’ estimated regression coefficients) and applied Bonferroni corrections to account for multiple comparisons. We conducted such pairwise comparisons only on pairs of treatments that are the subject of the hypotheses described above. In Section 4, we discuss all pairwise differences that were significant at the ¡ .05 level. f Finally, as a robustness check, we repeated the analysis described above while accounting for three different sets of covariates, by including them as additional independent variables:
Participant-level variables: we included seven variables concerning the evaluator: how often they use GenAI models for generating 1) images and 2) text; whether they had already participated in a study in which they were asked to 3) generate and 4) evaluate images; and whether they have any training in professions related to 5) art, 6) computer science, or 7) law.
@@
Our results are robust to all three groups of covariates. That is, pairwise comparisons between treatments remain qualitatively and quantitatively similar upon including any of the three groups of covariates, barring minor changes in the significance of borderline results. For simplicity and brevity, we do not discuss the results of models with participant- and image-level covariates in the paper. However, we discuss order effects when the coefficient is significant at the ¡ .05 level.
@@ 3.6. Data Collection & Participants
We discarded responses from 21 participants who failed any of two instructed response questions. Due to technical problems, responses from three participants had to be dropped because they were not saved completely. Some participants took part in the study more than once, in which case we only kept their first response. Finally, we discarded judgments made by (invested and uninvested) evaluators regarding images generated by creators that were removed due to attention check failures. Our final sample comprises 424 participants, out of which 95 were creators, 165 were invested evaluators, and 164 were uninvested evaluators. All participants were paid 3.50 USD for their participation regardless of their treatment condition to keep monetary incentives constant (approximately 11.50 USD per hour).
Half of participants identified themselves as women, 47.4% as men, and 2.12% as non-binary. Participants’ mean age was 40.5 years old ( = 14.0), with the youngest respondent being 20 years old and the oldest 81 years old. Our sample is slightly more diverse than the US population in terms of race, with 10.1% of participants describing themselves as Asian, 15.3% as Black or African American, and 61.6% as White. Only 7.5% of participants reported having prior training in professions related to law, while 24.1% stated that they had training in computer science-related professions and 26.2% in art-related occupations.
In addition to the participants recruited for the main study, we recruited an additional sample of evaluators for determining which art should be included in the exhibition. In the main study, only creators’ submissions received evaluations from other participants. Specifically, each of the creators’ submissions was evaluated by approximately 13 other participants from the pool of invested and uninvested evaluators. On the other hand, invested evaluators’ submissions were only evaluated by themselves. Since both creators and invested evaluators were eligible to receive awards based on how others rate their submissions, we recruited 115 participants to rate 20 of the invested evaluators’ submissions each, resulting in approximately 13 evaluations for each submission to the art show. These participants were paid 1.60 USD (approximately 11.30 USD per hour).
@@ 4. Results
A total of 260 participants submitted an image to the AI art exhibition (95 creators and 165 invested evaluators). These participants generated a median of two images (mean = 4.33, = 5.08). Although 90 of these participants generated only one image, 30 of them created 10 or more. Both creators and invested evaluators generated a median of two images, with creators creating more on average (mean = 5.01, = 6.40) than invested evaluators (mean = 3.93, = 4.1). Participants took an average of 5.86 minutes ( = 5.86) to create and select an image for the exhibition. Invested evaluators (mean = 5.94 minutes, = 5.66) spent slightly more time generating images than creators (mean = 5.72, = 6.23). Figure 3 presents some example images.
@@ 4.1. RQ1: Perceived Creativity, Effort, and Skills
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = 4.505 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = 2.503 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = 2.203 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = -2.844 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = 3.948 | |||
@@
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = 3.616 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = 1.752 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = 3.123 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = -2.284 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = 4.247 | |||
@@
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = 0.953 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = -1.111 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = -1.746 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = -2.803 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = 1.241 |
Figure 4A shows participants’ mean evaluations regarding the creativity, effort, and skills involved in generating images with GenAI. Participants somewhat agreed that creativity (mean = 0.656, = 1.72) and effort (mean = 0.488, = 1.79) were necessary to create the images. In contrast, judgments concerning skills were closer zero (mean = -0.070, = 1.82), meaning that on average participants neither agreed nor disagreed that creators used their skills to generate images with GenAI. Table 1 presents pairwise comparisons of perceived creativity, effort, and skills between treatment conditions. Below, we discuss results of pairwise comparisons for each factor separately.
@@ 4.1.1. Creativity
Concerning creativity, uninvested evaluators rated images lower than creators (diff = ) and invested evaluators, both when the latter evaluated their own creations (diff = ) and other people’s images (diff = ). That is, our results regarding creativity support H4, and they partially support H1. Finally, we found an effect in the opposite direction than the one hypothesized in H3—invested evaluators assigned higher creativity scores to creators than uninvested evaluators did.
@@ 4.1.2. Effort
Our analysis shows that uninvested evaluators assign lower effort ratings than creators (diff = ) and invested evaluators judging their own images (diff = ), offering partial support to H1 and H4. We also find that invested evaluators assign higher effort ratings to their own creations than to others’ (diff = ), in line with H2.
@@ 4.2. RQ2: Perceived Authorship
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = -2.126 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = -4.123 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = -6.386 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = -2.509 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = -1.326 | |||
@@
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = 1.355 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = 2.142 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = 1.762 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = 0.96 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = 0.011 | |||
@@
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = 0.592 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = -0.031 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = -1.695 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = -0.747 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = -0.119 | |||
@@
| Creators (Own) - Uninvested Evaluators (Others) - H1 & H4 | t(1515) = -0.162 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = 0.302 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = -0.067 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 & H4 | t(1515) = 0.557 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = -0.587 |
Figure 4B shows how participants attributed authorship between the user, the AI model, the company that developed the AI model, and data contributors. Participants somewhat agreed that users (mean = 0.537, = 1.69) and data contributors (mean = 0.571, = 1.69) are authors of AI-generated images. Judgments concerning the AI model were more uncertain (mean = 0.037, = 1.78), with evaluators on average neither agreeing nor disagreeing that the AI model itself is an author. The company that developed the AI model had the lowest perceived authorship (mean = -0.476, = 1.71). Table 2 presents pairwise comparisons of perceived authorship between treatment conditions. Below, we discuss results for each entity separately.

### v2

@@ 3.4. Hypotheses: Egocentricity, Competition, and Experience Effects
We build upon prior work studying egocentric effects (see Section 2.5) and study how they may emerge in the context of GenAI. We formed two hypotheses concerning egocentric effects on lay perceptions of AI-generated images in relation to copyright. In H1, we compare the judgments of different participants about the same image. Namely, we hypothesize that the creator of an image will judge their own image more favorably than other participants. In H2, we compare the judgments of the same participant about different images. We hypothesize that participants will judge images they create more favorably than they will judge images created by others.
@@
H1) Egocentric effect between participants: Creators will 1) evaluate images more favorably and 2) be more likely to identify users as authors and right-holders when judging their own creations compared to invested and uninvested evaluators judging the same images.
@@
H2) Egocentric effect between images: Invested evaluators will 1) evaluate images more favorably and 2) be more likely to identify users as authors and right-holders when judging their own creations than when judging other people’s images.
@@
We also explore the aforementioned competition effect. Creators’ and invested evaluators’ outcomes depended on other participants’ assessments since only the highest-rated images would receive monetary rewards. Invested evaluators had the opportunity to inflate their own relative rating not only by giving high ratings to their own submission (in line with H2), but also by giving low scores to other people’s submissions. That is, they could benefit from sabotaging their competition. In contrast, uninvested evaluators had no incentive to sabotage other participants since they were not competing for the monetary rewards. Therefore, we hypothesize that the incentive structure associated with our juried art exhibition may lead those with something at stake (invested evaluators) to judge others’ submissions more harshly than those who do not have anything to gain from doing so (uninvested evaluators).
@@
H3) Competition effect: Invested evaluators will 1) evaluate other people’s images less favorably and 2) be less likely to identify other users as authors and right-holders than uninvested evaluators.
@@
Finally, we also make hypotheses based on the experience effect. Uninvested evaluators did not have an opportunity to interact with the GenAI model and to create AI-generated art in our experiments. On the other hand, creators and invested evaluators repeatedly interacted with the model throughout the tutorial and creation task, observing how changing their inputs influences the model’s outputs and learning how to use it to create AI-generated art. Consequently, we hypothesize that the experience of interacting with the GenAI model will influence participants’ perceptions of AI-generated art.
@@
H4) Experience effect: Participants who interacted with the GenAI model (i.e., creators and invested evaluators) will 1) evaluate images and 2) attribute authorship and rights differently than those who did not use the GenAI model (i.e., uninvested evaluator).
@@
Finally, we emphasize that while we gather data on how participants attribute authorship and rights to various entities, our hypotheses focus on attributions of authorship and rights to one specific entity: users. We investigate the remaining entities exploratively.
@@ 3.5. Analysis Plan
We used regressions to analyze our data. We treated participants’ assessments of factors, authorship, and score evaluations as continuous dependent variables in linear regressions. Participants’ attributions of rights were treated as binary dependent variables and modeled using logit regressions. To account for repeated measurements across participants and images (i.e., participants evaluated several images, and images were evaluated by several participants), we initially planned to use mixed-effects linear regressions with crossed random intercepts for images and participants. However, due to convergence issues, we instead opted for regressions with two-way clustered standard errors (Correia et al., 2016).
Our primary independent variable is a dummy variable encoding both the i) treatment condition to which the participants were assigned and ii) whether the data point refers to an evaluation of their own image or someone else’s creation. Hence, our dummy variable has four levels representing 1) creators judging their own images; 2) invested evaluators rating their own creations; 3) invested evaluators assessing others’ images; and 4) uninvested evaluators judging others’ submissions.
We tested for differences between pairs of treatments by estimating their contrasts (i.e., by estimating the difference between the treatments’ estimated regression coefficients) and applied Bonferroni corrections to account for multiple comparisons. We conducted such pairwise comparisons only on pairs of treatments that are the subject of the hypotheses described above. In Section 4, we discuss all pairwise differences that were significant at the ¡ .05 level.
Finally, as a robustness check, we repeated the analysis described above while accounting for three different sets of covariates by including them as additional independent variables:
Participant-level variables: we included ten variables concerning the evaluator: their self-reported 1) gender, 2) age, and 3) race; how often they use GenAI models for generating 4) images and 5) text; whether they had already participated in a study in which they were asked to 6) generate and 7) evaluate images; and whether they have any training in professions related to 8) art, 9) computer science, or 10) law.
@@
Our results are robust to all three groups of covariates. That is, pairwise comparisons between treatments remain qualitatively and quantitatively similar upon including any of the three groups of covariates, barring minor changes in the significance of borderline results. For simplicity and brevity, we do not discuss the results of models with participant- and image-level covariates in the paper. However, we discuss order effects when the coefficient is significant at the ¡ .05 level. We present all regressions—with and without covariates—in the Appendix and make all of our data available for replication.
@@ 3.6. Data Collection & Participants
Due to technical problems, responses from two participants had to be dropped because they were not saved completely. We discarded responses from 16 participants who failed any of two instructed response questions. Finally, we discarded judgments made by (invested and uninvested) evaluators regarding images generated by creators that were removed due to attention check failures. Our final sample comprises 432 participants, out of which 95 were creators, 169 were invested evaluators, and 168 were uninvested evaluators. All participants were paid $3.50 USD for their participation regardless of their treatment condition to keep monetary incentives constant (approximately $12.00 USD per hour).
Nearly half (49.77%) of participants identified themselves as women, 46.99% as men, and 3.23% as non-binary or chose not to disclose their gender. Participants’ mean age was 40.3 years old ( = 14.0), with the youngest respondent being 20 years old and the oldest 81 years old. Our sample is slightly more diverse than the US population in terms of race, with 10.19% of participants describing themselves as Asian, 15.05% as Black or African American, and 61.34% as White. Only 7.41% of participants reported having prior training in professions related to law, while 23.61% stated that they had training in computer science-related professions and 26.39% in art-related occupations.
In addition to the participants recruited for the main study, we recruited an additional sample of evaluators to determine which art should be included in the exhibition. In the main study, only creators’ submissions received evaluations from other participants. Specifically, each of the creators’ submissions was evaluated by approximately 13 other participants from the pool of invested and uninvested evaluators. On the other hand, invested evaluators’ submissions were only evaluated by themselves. Since both creators and invested evaluators were eligible to receive awards, we recruited 115 additional participants to rate 20 of the invested evaluators’ submissions each, resulting in approximately 13 evaluations for each submission to the art show. Participants in this additional study only evaluated images with respect to quality to determine awards. These participants were paid 1.60 USD (approximately $12.00 USD per hour).
@@ 4. Results
Six example images. The top-left image depicts a person with dark skin and braids in front of a flower field. The top-center image depicts an Asian woman holding a baby in a fairytale-like setting. The top-right image portrays a forest. The bottom-left image is an abstract picture with the symbol of the female gender symbol in the center. The bottom-center image depicts a dark universe with some stars. The bottom-right image is a colorful abstract image with circular patterns.
A total of 264 participants submitted an image to the AI art exhibition (95 creators and 169 invested evaluators). These participants generated a median of two images (mean = 4.35, = 5.10). Although 90 of these participants generated only one image, 31 of them created 10 or more. Creators generated a median of 2 images (mean = 5.01, = 6.40), while invested evaluators created a median of 3 images (mean = 3.98, = 4.16). Participants took an average of 5.87 minutes ( = 5.86) to create and select an image for the exhibition. Invested evaluators (mean = 5.95 minutes, = 5.61) spent slightly more time generating images than creators (mean = 5.73, = 6.23). Figure 3 presents some example images.
@@ 4.1. RQ1: Perceived Creativity, Effort, and Skills
The figure consists of two panels. Panel A (on the left) is a bar graph representing the mean values of perceived creativity, effort, and skills. The plot also includes mean values depending on the treatment condition. Panel B (on the right) is an equivalent bar graph that illustrates mean attributions of authorship to the user, the AI model, the company that developed the AI model, and data contributors. The plot also shows mean values depending on the treatment condition.
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = 4.704 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = 2.431 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = 2.179 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = -3.226 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = 4.250 | |||
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = 3.899 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = 1.809 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = 3.252 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = -2.556 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = 4.512 | |||
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = 1.159 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = -1.097 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = -1.592 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = -3.055 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = 1.576 |
Figure 4A shows participants’ mean evaluations regarding the creativity, effort, and skills involved in generating images with GenAI. Participants somewhat agreed that creativity (mean = 0.642, = 1.73) and effort (mean = 0.458, = 1.81) were necessary to create the images. In contrast, judgments concerning skills were closer to zero (mean = -0.087, = 1.82), meaning that, on average, participants neither agreed nor disagreed that creators used their skills to generate images with GenAI. Table 1 presents pairwise comparisons of perceived creativity, effort, and skills between treatment conditions. Below, we discuss the results of pairwise comparisons for each factor separately.
@@ 4.1.1. Creativity
Concerning creativity, uninvested evaluators rated images lower than creators (diff = ) and invested evaluators, both when the latter evaluated their own creations (diff = ) and other people’s images (diff = ). That is, our results regarding creativity support H4, and they partially support H1. We also found an effect in the opposite direction than the one hypothesized in H3—invested evaluators assigned higher creativity scores to creators than uninvested evaluators did.
@@ 4.1.2. Effort
Our analysis shows that uninvested evaluators assigned lower effort ratings than creators (diff = ) and invested evaluators judging their own images (diff = ), offering partial support to H1 and H4. We also find that invested evaluators assigned higher effort ratings to their own creations than to others’ (diff = ), in line with H2.
@@ 4.2. RQ2: Perceived Authorship
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = -1.855 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = -4.018 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = -6.300 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = -2.728 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = -1.017 | |||
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = 1.442 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = 2.132 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = 1.788 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = 0.846 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = 0.133 | |||
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = 0.619 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = -0.015 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = -1.557 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = -0.764 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = -0.012 | |||
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = -0.208 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = 0.238 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = -0.271 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = 0.542 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = -0.685 |
Figure 4B shows how participants attributed authorship between the user, the AI model, the company that developed the AI model, and data contributors. Participants somewhat agreed that users (mean = 0.508, = 1.71) and data contributors (mean = 0.579, = 1.70) are authors of AI-generated images. Judgments concerning the AI model were more uncertain (mean = 0.031, = 1.79), with evaluators, on average, neither agreeing nor disagreeing that the AI model itself is an author. The company that developed the AI model had the lowest perceived authorship (mean = -0.479, = 1.72). Table 2 presents pairwise comparisons of perceived authorship between treatment conditions. We discuss the results for each entity separately.